# No Detail Left Behind: Revisiting Self-Retrieval for Fine-Grained Image Captioning

### v1

@@ 3.3 Experiment 2: Benchmarking Visual Caption Boosting with Self-Retrieval
We evaluate the quality of captions adopted in VCB in Table 2. Along with TrueMatch, we also adopt the SR evaluation strategy of Dessì et al. (2023) with 100 random distractors (RD100). On RD100, BlendCap outperforms original COCO captions by a large margin of 8% on R@1 and 6.2% on ClipScore, confirming that human annotations capture complementary visual aspects of the same image. However, this sizeable gap shrinks to 2.5% on TrueMatch (bags of size #3). This confirms that even though annotated captions blended together work better, they inherently lack fine-grained details (see Figure 2) necessary to perform well on TrueMatch. HolisticCap, on the other hand, yields remarkable performance gains over COCO and BlendCap across both setups: RD100 and all bag sizes of TrueMatch. This demonstrates the effectiveness of VCB in instilling fine-grained information into standard image captioning datasets.
Effectiveness of anchoring visual captions. Table 2 shows that the standalone visual caption (generated using InstructBLIP) is slightly worse than BlendCap. However, creating HolisticCap by anchoring the visual caption in human annotations results in about +4% improvement over BlendCap on TrueMatch. By drawing comparisons to Table 1, we observe that HolisticCap, which uses InstructBLIP, achieves performance on par to larger (LLaVa 1.6) and newer (Cambrian-1-3B) MLLMs.
@@
In summary, SR within TrueMatch enables fine-grained evaluation of captioning systems, probing different facets of visual discrimination. It encourages captions be to unique while also capturing semantic similarity with images, addressing the limitations of traditional captioning metrics.
@@ 4.1 Methodology
Similar to Dessì et al. (2023), we adopt ClipCap (Mokady et al., 2021), a lightweight simplification of the modern MLLMs (e.g. LLaVA, InstructBLIP, Cambrian). ClipCap connects a pretrained visual encoder (CLIP (Radford et al., 2021)) to a pretrained language model (GPT-2 (Radford et al., 2019)) through a simple MLP adapter. The adapter is tasked with mapping the rich visual embeddings from CLIP into a fixed number of prefix tokens. These tokens capture essential visual information and guide the language model towards generating an image-conditioned caption.
@@
models the training data distribution. Specifically, captioning models are often teacher-forced to learn the word (token) distribution that maximizes the log-likelihood of the ground-truth captions given an input image. However, this results in strong language modeling priors resulting in generic captions that use a small vocabulary (see Table 6).
@@ 4.2 Experiment Setup and Metrics
We evaluate the quality of captions with NLG metrics: BLEU-4 and CIDEr. On RD100, a setup with 100 random distractors (Dessì et al., 2023), we use R@1 and ClipScore. On TrueMatch, we report R@1.
@@ 4.3 Experiment 3: Improving SR with Richer MLE Initialization
Following the findings of Section 3.3, we use TrueMatch to investigate the impact of different MLE initializations on SR fine-tuning in Table 3. We study SR in the same setting as previous works (Dessì et al., 2023), i.e. we only fine-tune the language model (SR-L).
Table 3: Impact of MLE initialization and fine-tuning with SR. Results across different training approaches (MLE, SR) and captions (COCO, BlendCap, HolisticCap) show the effectiveness of improved initialization with HolisticCap and SR training. As CLIP is kept frozen, we refer to this as SR-L. Row 1 is ClipCap (Mokady et al., 2021) and row 4 is Discritune (Dessì et al., 2023).
MLE Training NLG metrics RD100 TrueMatch
Dataset B C R@1 ClipSc #3 #5 #7
1 COCO 33.2 107.8 76.5 29.9 50.8 36.5 33.8
2 BlendCap MLE 19.7 80.2 84.9 31.6 51.8 42.5 42.2
3 HolisticCap 16.3 26.6 84.7 31.7 55.9 43.7 42.2
4 COCO 32.3 108.2 83.7 30.2 53.3 42.3 38.4
5 BlendCap SR-L 19.6 80.1 87.5 31.7 54.7 47.5 44.2
6 HolisticCap 17.0 27.8 87.6 31.9 59.1 47.7 46.5
@@
Table 3 rows 1-3 report the performance of MLE training on captioning datasets with increasing level of detail. We observe that improving the training captions results in a significant performance boost. Training with HolisticCap beats COCO by +8.2% on RD100 R@1 and +5.1% on TrueMatch #3. However, comparing performance against the ground-truth HolisticCap directly (from Table 2), we see that MLE training on HolisticCap yields -6.6% on RD100 R@1 and -1.6% on TrueMatch #3. This indicates that even though HolisticCap improves performance, MLE training tends to generate generic and non-discriminative descriptions.
VCB breaks CIDEr evaluation.
We see a significant drop in CIDEr scores for BlendCap and HolisticCap compared to COCO in Table 3. Upon further analysis, we believe that CIDEr rewards brevity and struggles to effectively evaluate with reference sets containing longer descriptive captions. For example with HolisticCap as ground-truth, Section C.6 shows some valid captions that unfortunately obtain scores close to 0.
@@
SR-L overcomes the limitations of MLE by making the captions more discriminant, leading to consistent improvements in RD100 and TrueMatch across all 3 captions (compare rows 1-4, 2-5, 3-6). Providing a richer initialization to SR-L through VCB enhances fine-grained visual discrimination. Row 6 outperforms DiscriTune (row 4) with +3.9% RD100 R@1 and +5.8% TrueMatch #3. This illustrates the importance of a good MLE initialization for SR and the effectiveness of VCB.
@@
Interestingly, SR-L fine-tuning narrows this gap dramatically from +8.4% to a mere +3.8% (rows 4, 5). This indicates that SR fine-tuning teaches the captioner to uncover semantic information within its embedding space, even when it was initially obscured by MLE. Furthermore, the mostly unchanged BLEU-4 and CIDEr scores hint that the captioner remains faithful to the GT captions for both COCO (rows 1, 4) and BlendCap (rows 2, 5). This is remarkable, as it demonstrates that SR fine-tuning unlocks latent semantic information and steers the language model while preserving the initial data distribution.
@@
Self-retrieval guides the captioner away from its language modeling priors by unlocking latent semantic information and steering the language model to generate discriminative captions.
4.4 Trade-off plaguing SR fine-tuning: Retrieval Performance vs Faithfulness
While SR fine-tuning removes language modeling priors from captioners by making them more discriminant, we observe that they have a tendency to become less faithful to the GT captions upon extended training. We contrast these two tendencies of SR fine-tuning in Figure 3, comparing CIDEr and RD100 R@1 scores.
@@ 4.4 Trade-off plaguing SR fine-tuning: Retrieval Performance vs Faithfulness
While SR performance continually improves (+7.9%) with extended training, a corresponding degradation in CIDEr (-15.4%) is observed. This trade-off reveals that SR has the capacity to elicit better retrieval from the captioning system, even at the cost of generating lower-quality captions. We further investigate with qualitative examples in Section C.7. Dessì et al. (2023) attribute this deterioration to the model diverging from the GT distribution during SR-L fine-tuning and argue that it is desirable. While this is partially true (see Figure 7), our investigation also reveals that captioners trained with vanilla SR-L, in a dash to enhance retrieval performance, build a propensity to hallucinate (see Figure 8) resulting in poor CIDEr scores. This tendency is only exacerbated with extended training due to further degradation in NLG metrics. These findings reveal an important bottleneck in SR fine-tuning:
@@
Captioners fine-tuned with SR-L suffer from a trade-off between retrieval performance and caption faithfulness, showing a tendency to hallucinate details and deviate from ground-truth captions.
Finally, the continued improvement in RD100 R@1 over 100 epochs indicates that vanilla SR-L fine-tuning fails to saturate SR’s capacity to instill fine-grained visual discrimination in captioning systems. This motivates us to explore efficient SR fine-tuning strategies while mitigating their propensity to hallucinate. Since a higher learning rate destabilizes the policy gradients, we turn to: (1) fine-tuning the visual encoder (CLIP) with SR; and (2) adopting a curriculum over bags of hard negatives.
@@ 4.5 Experiment 4: Fine-tuning CLIP with Self-Retrieval
Further qualitative analysis of the captions generated by SR-L and SR-V in Figure 9, reveal that captioners fine-tuned with SR-V struggle to bind attributes to the correct objects. For instance, while SR-V captures the red color of the crane, it binds it with the incorrect object (train). This is likely due to two reasons: (1) the retrieval objective treats the caption as a bag of words and does not promote compositionality; and (2) CLIP has been shown to struggle with attribute binding (Lewis et al., 2022, Yuksekgonul et al., 2023, Hsieh et al., 2024). We also observe that SR-V exacerbates SR’s tendency to hallucinate (see Figure 9).
@@
On TrueMatch, while SR-V fine-tuning provides substantial improvements over SR-L for COCO and BlendCap (rows 1-2, 4-5), we observe relatively smaller gains for HolisticCap (rows 7-8). This indicates that the vanilla SR-L objective of retrieving against 99 random distractors is not challenging enough to provide a good learning signal for models pretrained on fine-grained HolisticCap. While SR-LV does better than SR-V on HolisticCap (rows 8-9), the gap between BlendCap (row 6) and HolisticCap (row 9) still remains small. To address this, we devise a curriculum with hard negatives, discussed in the next section, to increase the difficulty of SR fine-tuning objective.
@@
Fine-tuning the visual module with self-retrieval instills fine-grained visual details in the captioning system but struggles with attribute binding and increased hallucinations.
@@ 4.6 Experiment 5: Self-Retrieval Fine-tuning with Bags and Curriculum
Table 5: Ablation of fine-tuning with the bag curriculum (BagCurri) as compared against vanilla SR (SR-L). Row 1 is Dessì et al. (2023). (Green) numbers indicate absolute improvement and (red), a degradation.
@@
6 Holistic- SR-L 31.3 (3.5) 88.5 (0.9) 58.4 (0.7) 50.0 (2.3) 49.5 (3.0)
7 Cap SR-V 25.8 (2.0) 90.6 (3.0) 64.0 (4.9) 54.6 (6.9) 54.1 (7.6)
8 SR-LV 29.7 (1.9) 91.3 (3.7) 65.2 (6.1) 57.1 (9.4) 57.1 (10.6)
The SR experiments above used the same strategy as Dessì et al. (2023) with 99 random distractors. Instead, in this experiment, we fine-tune with bags of highly similar images within a training minibatch (see Section B.2). Retrieving against multiple hard negatives flattens the softmax distribution resulting in a stronger learning signal. We also propose a curriculum over bag sizes (BagCurri) to more optimally leverage the contrastive reward (see Section B.3). SR fine-tuning with BagCurri instills fine-grained visual details in the model, essential for discriminating highly similar images.
@@
We evaluate the impact of fine-tuning different components of the captioning system (CLIP and GPT-2) with BagCurri compared to vanilla SR-L in Table 5. We find that SR-V fine-tuning with our curriculum forces CLIP to learn fine-grained visual features resulting in a substantial +5% to +11% improvement over vanilla SR-L (rows 3 vs. 1, 7 vs. 5) on TrueMatch. Interestingly, SR-V with BagCurri renders COCO (row 3) comparable to HolisticCap (row 7) on TrueMatch #3. However, these disproportionate gains for COCO come at the expense of CIDEr scores, which plummet -22.4% compared to SR-L. Further qualitative analysis of the generated captions shows that SR-V fine-tuning with BagCurri amplifies the trends observed in Section 4.5 – while it boosts retrieval performance, it exacerbates the inability of the captioner to bind attributes to the correct objects and increases hallucinations (see Figure 9). In contrast, BagCurri may be too challenging for SR-L fine-tuning, as evident by the small improvements of +0% to +4% over vanilla SR-L (rows 2 vs. 1, 6 vs. 5) on RD100 or TrueMatch. However, similar to the findings of Section 4.3, we find that fine-tuning SR-L with BagCurri benefits heavily from the rich initialization provided by VCB. While the CIDEr score drops for COCO when using SR-L with BagCurri (-4.8%, row 2), for HolisticCap we observe an improvement (+3.5%, row 6).
@@
Above findings indicates that BagCurri has the capacity to enhance both: (1) visual discrimination ability of the captioner with SR-V; and (2) adherence to GT captions with SR-L when provided with a rich initialization. Motivated by this, we initialize the captioner with HolisticCap and fine-tune both the language and visual components (SR-LV) with BagCurri (row 8), retaining the best of both worlds. SR-LV+BagCurri further enhances SR-V’s already impressive performance on TrueMatch (achieving +6-10% over vanilla SR-L) while also improving the CIDEr score over vanilla SR-L by +1.9% (rows 5, 8). This is momentous, as without modifying the model’s architecture or reward function, we outperform vanilla SR-L (Dessì et al., 2023) by an average +15% across all bag sizes on TrueMatch (row 1, 8), while circumventing the caption faithfulness trade-off plaguing current SR approaches.
@@
Providing a rich initialization with VCB and fine-tuning SR-LV with BagCurri instills fine-grained visual details in the captioner while preserving its faithfulness to ground-truth captions.
4.7 Impact on Generating Diverse Caption

### v2

@@ 3.3 Experiment 2: Benchmarking Visual Caption Boosting with Self-Retrieval
We evaluate the quality of captions adopted in VCB in Table 2. Along with TrueMatch, we also adopt the SR evaluation strategy of Dessì et al. (2023) with 99 random distractors (RD100). On RD100, BlendCap outperforms original COCO captions by a large margin of 8% on R@1 and 6.2% on ClipScore, confirming that human annotations capture complementary visual aspects of the same image. However, this sizeable gap shrinks to 2.5% on TrueMatch (bags of size #3). This confirms that even though annotated captions blended together work better, they inherently lack fine-grained details (see Figure 2) necessary to perform well on TrueMatch. HolisticCap, on the other hand, yields remarkable performance gains over COCO and BlendCap across both setups: RD100 and all bag sizes of TrueMatch. This demonstrates the effectiveness of VCB in instilling fine-grained information into standard image captioning datasets.
We study the effectiveness of anchoring visual captions in producing more succinct (see Section A.3) and fine-grained captions. Table 2 shows that standalone VisualCap (generated using the MLLM) is slightly worse than BlendCap. However, anchoring VisualCap in human annotations to create HolisticCap results in +4% improvement over BlendCap on TrueMatch. Additionally, by drawing comparisons to Table 1, we observe that HolisticCap achieves performance on par with larger (LLaVa 1.6) and newer (Cambrian-1-3B) MLLMs compared to InstructBLIP.
@@
In summary, SR within TrueMatch enables fine-grained evaluation of captioning systems, probing different facets of visual discrimination. It requires captions to be unique even for semantically similar images, thereby addressing the limitations of traditional captioning metrics.
@@ 4.1 Methodology
Similar to Dessì et al. (2023), we adopt ClipCap (Mokady et al., 2021), a lightweight simplification of modern MLLMs (e.g. LLaVA, InstructBLIP, Cambrian). ClipCap connects a pretrained visual encoder (CLIP (Radford et al., 2021)) to a pretrained language model (GPT-2 (Radford et al., 2019)) through a simple MLP adapter. The adapter is tasked with mapping the rich visual embeddings from CLIP into a fixed number of prefix tokens. These tokens capture essential visual information and guide the language model towards generating an image-conditioned caption.
@@
models the training data distribution. Specifically, captioning models are often teacher-forced to learn the word (token) distribution that maximizes the log-likelihood of the ground-truth captions given an input image. However, this results in strong language modeling priors resulting in generic captions that use a small vocabulary (see Section 4.7).
@@ 4.2 Experiment Setup and Metrics
We evaluate the quality of captions with NLG metrics: BLEU-4 and CIDEr, keeping the reference set consistent with the training distribution. On RD100, a setup with 99 random distractors (Dessì et al., 2023), we use R@1 and ClipScore. On TrueMatch, we report R@1.
@@ 4.3 Experiment 3: Improving SR with Richer MLE Initialization
Following the findings of Section 3.3, we use TrueMatch to investigate the impact of different MLE initializations on SR fine-tuning in Table 3. We study SR in the same setting as previous work (Dessì et al., 2023): (1) we only fine-tune the language model (SR-L), but using LoRA adapters; and (2) until Section 4.6 we use 99 random distractors to fine-tune with the SR reward.
Table 3: Results for different training methods and captions show the effectiveness of improved initialization with HolisticCap and SR-L.
MLE Training CIDEr RD100 TrueMatch
Dataset R@1 ClipSc #3 #5 #7
1 COCO 107.8 76.5 29.9 50.8 36.5 33.8
2 BlendCap MLE 80.2 84.9 31.6 51.8 42.5 42.2
3 HolisticCap 26.6 84.7 31.7 55.9 43.7 42.2
4 COCO 108.2 83.7 30.2 53.3 42.3 38.4
5 BlendCap SR-L 80.1 87.5 31.7 54.7 47.5 44.2
6 HolisticCap 27.8 87.6 31.9 59.1 47.7 46.5
@@
Table 3 rows 1-3 report the performance of MLE training on captioning datasets with increasing level of detail. We observe that improving the training captions results in a significant performance boost. Training with HolisticCap beats COCO by +8.2% on RD100 R@1 and +5.1% on TrueMatch #3. However, comparing performance against the ground-truth HolisticCap directly (from Table 2), we see that MLE training on HolisticCap yields -6.6% on RD100 R@1 and -1.6% on TrueMatch #3. This indicates that even though HolisticCap improves performance, MLE training (i) generates generic descriptions that may not be aligned with the image content, and (ii) leads to object hallucinations (see Section C.7).
@@
SR-L reduces MLE’s preference towards statistically probable phrases resulting in more discriminant captions, reduced object hallucinations and consistent improvements in RD100 and TrueMatch across datasets (compare rows 1-4, 2-5, 3-6). Providing a richer initialization to SR-L through VCB enhances fine-grained visual discrimination. Row 6 outperforms DiscriTune (row 4) (Dessì et al., 2023) with +3.9% RD100 and +5.8% TrueMatch #3. This illustrates the importance of a good MLE initialization for SR and the effectiveness of VCB.
Longer reference sets hurt CIDEr.
We see a significant drop in CIDEr scores for BlendCap and HolisticCap compared to COCO in Table 3. Upon further analysis, we believe that CIDEr rewards brevity. Section C.6 shows some valid short captions that unfortunately obtain scores close to 0.
@@
Interestingly, SR-L fine-tuning narrows this gap dramatically from +8.4% to a mere +3.8% (rows 4, 5). This indicates that SR fine-tuning teaches the captioner to uncover semantic information within its embedding space, even when it was initially obscured by MLE. Furthermore, the mostly unchanged CIDEr scores hint that the captioner remains faithful to the GT captions for both COCO (rows 1, 4) and BlendCap (rows 2, 5). This is remarkable, as it demonstrates that SR fine-tuning unlocks latent semantic information and steers the language model while preserving the initial data distribution.
@@
Self-retrieval steers the language model to generate discriminative captions by uncovering latent semantic information that was initially obscured by MLE.
4.4 Trade-off plaguing SR fine-tuning: Retrieval Performance vs. Caption Faithfulness
While SR fine-tuning removes language modeling priors from captioners by making them more discriminant, we observe that they have a tendency to become less faithful to the GT captions upon extended training. We contrast these two tendencies of SR fine-tuning in Figure 3, comparing CIDEr and RD100 R@1 scores. We train for 100 epochs compared to the default 20 epochs used for all other experiments.
@@
While SR performance continually improves (+7.9%) with extended training, a corresponding degradation in CIDEr (-15.4%) is observed. This trade-off reveals that SR has the capacity to elicit better retrieval from the captioning system, even at the cost of generating lower-quality captions. We further investigate with qualitative examples in Section C.8. Dessì et al. (2023) attribute this deterioration to the model diverging from the GT distribution during SR-L fine-tuning and argue that it is desirable. While this is partially true (see Figure 7), our investigation also reveals that captioners trained with vanilla SR-L, in a dash to enhance retrieval performance, build a propensity to hallucinate attributes (see Figure 8) resulting in poor CIDEr scores. This tendency is only exacerbated with extended training resulting in further degradation of NLG metrics and increased object hallucinations (see Section C.7). These findings reveal an important bottleneck in SR fine-tuning:
@@
Captioners fine-tuned with SR-L suffer from a trade-off between retrieval performance and faithfulness, often hallucinating details and deviating from GT captions with extended training.
Finally, the continued improvement in RD100 R@1 over 100 epochs indicates that vanilla SR-L fine-tuning fails to saturate SR’s capacity to instill fine-grained visual discrimination in captioning systems. This motivates us to explore better SR fine-tuning strategies while mitigating their propensity to hallucinate. Since a higher learning rate destabilizes the policy gradients, we turn to: (1) fine-tuning the visual encoder (CLIP) with SR; and (2) adopting a curriculum over bags of hard negatives.
@@ 4.5 Experiment 4: Fine-tuning CLIP with Self-Retrieval
Further qualitative analysis of the captions generated by SR-L and SR-V in Figure 9, reveal that captioners fine-tuned with SR-V struggle to bind attributes to the correct objects. For instance, while SR-V captures the red color of the forklift, it binds it with the incorrect object (train). This is likely due to two reasons: (1) the retrieval objective treats the caption as a bag of words and does not promote compositionality; and (2) CLIP has been shown to struggle with attribute binding (Lewis et al., 2022, Hsieh et al., 2024).
@@
On TrueMatch, while SR-V fine-tuning provides substantial improvements over SR-L for COCO and BlendCap (rows 1-2, 4-5), we observe relatively smaller gains for HolisticCap (rows 7-8). This indicates that the vanilla SR-L objective of retrieving against 99 random distractors (RD100) is not challenging enough to provide a good learning signal for models pretrained on fine-grained HolisticCap. While SR-LV does better than SR-V on HolisticCap (rows 8-9), the gap between BlendCap (row 6) and HolisticCap (row 9) still remains small. To address this, we devise a curriculum with hard negatives, discussed in the next section, to increase the difficulty of SR fine-tuning objective.
@@
Fine-tuning the visual module with self-retrieval instills fine-grained visual details in the captioning system, while increasing hallucinations and struggling with attribute binding.
@@ 4.6 Experiment 5: Self-Retrieval Fine-tuning with Bags and Curriculum
Table 5: Ablation of fine-tuning with the bag curriculum (BagCurri) as compared against vanilla SR (SR-L). Row 1 is Dessì et al. (2023).
@@
6 Holistic- SR-L 31.3 88.5 58.4 50.0 49.5
7 Cap SR-V 25.8 90.6 64.0 54.6 54.1
8 SR-LV 29.7 91.3 65.2 57.1 57.1
The above SR experiments were trained using 99 random distractors (Dessì et al., 2023). Instead, in this experiment, we fine-tune with bags of highly similar images within a training minibatch (see Section B.2). Retrieving against multiple hard negatives flattens the softmax distribution resulting in a stronger learning signal. We also propose a curriculum over bag sizes (BagCurri) to more optimally leverage the contrastive reward (see Section B.3). SR fine-tuning with BagCurri instills fine-grained visual details in the model, essential for discriminating highly similar images.
@@
We evaluate the impact of fine-tuning different components of the captioning system (CLIP and GPT-2) with BagCurri compared to vanilla SR-L in Table 5. We find that SR-V fine-tuning with our curriculum forces CLIP to learn fine-grained visual features resulting in a substantial +5% to +11% improvement over vanilla SR-L (rows 3 vs. 1, 7 vs. 5) on TrueMatch. Interestingly, SR-V with BagCurri renders COCO (row 3) comparable to HolisticCap (row 7) on TrueMatch #3. However, these disproportionate gains for COCO come at the expense of CIDEr scores, which plummet -22.4% compared to SR-L. Further qualitative analysis of the generated captions shows that SR-V fine-tuning with BagCurri amplifies the trends observed in Section 4.5 – while it boosts retrieval performance, it exacerbates the inability of the captioner to bind attributes to the correct objects and increases hallucinations (see Figure 9). In contrast, BagCurri may be too challenging for SR-L fine-tuning, as evident by the small improvements of +0% to +4% over vanilla SR-L (rows 2 vs. 1, 6 vs. 5) on RD100 or TrueMatch. Nevertheless, similar to the findings of Section 4.3, we find that fine-tuning SR-L with BagCurri benefits heavily from the rich initialization provided by VCB. While the CIDEr score drops for COCO when using SR-L with BagCurri (-4.8%, row 2 vs. 1), for HolisticCap we observe an improvement (+3.5%, row 6 vs. 5).
@@
The above findings indicate that BagCurri has the capacity to enhance both: (1) visual discrimination ability of the captioner with SR-V; and (2) adherence to GT captions with SR-L when provided with a rich initialization. Motivated by this, we initialize the captioner with HolisticCap and fine-tune both the language and visual components (SR-LV) with BagCurri (row 8), retaining the best of both worlds. SR-LV+BagCurri further enhances SR-V’s already impressive performance on TrueMatch (achieving +6-10% over vanilla SR-L) while also improving the CIDEr score over vanilla SR-L by +1.9% (rows 5, 8). This is momentous, as without modifying the model’s architecture or reward function, we outperform vanilla SR-L (Dessì et al., 2023) by an average +15% across all bag sizes on TrueMatch (row 1, 8), while circumventing the caption faithfulness trade-off plaguing current SR approaches.
@@
Providing a rich initialization with VCB and SR-LV fine-tuning with BagCurri instills fine-grained visual details in the captioner while preserving its faithfulness to GT captions.
4.7 Generating Diverse Captions